What is the distance between the points (3 , -6) and (-3 , 2)?

Grade:10

2 Answers

sahil
6 years ago
the distance can be calculated by distance formula that is [{x(1)-x(2)}^2+{y(1)-y(2)}^2]^1/2 now here {x(1),y(1)} are (3,-6) and {x(2),y(2)} are (-3,2) so putting we get [6^2+(-8)^2]^1/2=[36+64]^1/2=[100]^1/2=10
aditya kulkarni
6 years ago
to find the distance between the two points simply use the distance formula-
distance between points M(a.b) and N (x,y)  is given by ((x-a)2+(y-b)2)1/2
here in the above qurstion a=3, b=-6, x=-3, y=2
therefore by putting the values in the formula
we det distance=10 units
 
if you agree with my solution please click “Approve” below